Consider the following calculation:

x=(b−c) / a

The following experimental values and their errors are obtained:

a = 26.9 , Δa = 4.9

b = 46.1 , Δb = 2.6

c = 24 , Δc = 3.9

What is the absolute error in x (Δx)?

Note we are not asking for the value of x itself, only its error.
Answer #1
Relation Equation limit error
Subtraction Z = x - y ܙܠ+rܠ - 2ܠ
division Z =x/y AZ = 141 ( + Sy

so, error in subtraction

ܐܠܐܠ = ܠ = 2.6 + 3.9 = 6.5

Z = b - c = 46.1 - 24 = 22.1

now,

\small \Delta a = 3.9  

now, ST    = | 6.5 4.9、 「22126.9 = ( 0.294 + 0.182) = 0.476

now , X = \small \frac{b - c}{a}    = 46.1-24 269 = \small \frac{22.1}{26.9} =   0.8216

now, absolute error in the final calculation

|AX] = |x|(2+ day = (0.8216 \small \times 0.476 )=  0.391

hence absolute error (|\small \Delta X| ) = 0.391